A Bibliography of Publications of Axel Ruhe (1942–2015)

Axel Ruhe Mathematics Department University of California Berkeley CA 94720-3840

USA

Tel: +1 510 642 8570 FAX: +1 510 642 8204

E-mail: ruhe@math.berkeley.edu (Internet)

01 December 2023 Version 2.06

Abstract

This bibliography records publications of Axel Ruhe.

Title word cross-reference

QR [Ruh66a, Ruh66b].

-method [Ruh66a, Ruh66b].

1 [Ruh12d, Ruh13a]. 1982 [KR83]. 1985 [MC87]. 1993 [Ruh94a].

2 [Ruh13b, Ruh14b].

3 [Ruh13c].

4 [Ruh13d].

54-3 [Ruh14a].

70th [GKRV90]. 75th [Ruh94a].

Accelerated [Ruh79c]. advances [GGL94]. after [Ruh69]. Algebra [BDD^{+}95, Ruh82]. algebraic [BDDR00, Ruh70a]. Algol [Ruh66a]. Algorithm [KR80a, KR80b, Ruh70c, Ruh80d, Ruh98, ER82, HR90, Ruh80a, Ruh83b, Ruh93, Ruh94b, Ruh00b, Ruh00a]. Algorithms [ER87, MC87, Ruh70a, Ruh73, Ruh79c, Ruh79a, RW80, RS98, Ruh74a, Ruh75a, Ruh94c, Ruh94d, MC87]. almost [Ruh75b]. APL [Ruh80a]. applied [HR90]. approach [WRWD84]. approximation [MC87]. Arbitrary [Ruh68]. Arnoldi [Ruh83b, Ruh96]. Aspects [Ruh79a, Ruh83a]. Axel [BE15].

Band [Ruh79a]. Based [Ruh80d, MC87, Ruh75a, Ruh80a]. Between [Ruh80d]. Birthday [GKRV90, Ruh94a]. BIT [Ruh12d, Ruh13a, Ruh13b, Ruh13c, Ruh13d, Ruh14b, Ruh14c].

BIT50 [Ruh11a]. Book [Ruh95]. Bounds [Ruh70d]. branch [ZR86].

Carl [Ruh94a]. Carl-Erik [Ruh94a]. chemistry [AHR87]. closeness [Ruh75b]. Closest [Ruh87]. Collaborators [Ruh09a]. College [MC87]. collinearity [WRWD84]. Complex [AHR87, KR80a, KR80b, Ruh66a, Ruh66b, Ruh10e, Ruh94b]. components [Ruh74b]. Computation [KR80a, KR80b, Ruh77, RK78, Ruh79a, Ruh84, RS98, OR06, Ruh74b]. Computations [Ruh82, Ruh95]. Computing [Ruh96]. conditioned [Ruh70b]. Conference [KR83, MC87]. conjugate [RW72]. constants [AHR87]. constraints [Ruh79b]. Contents [Ruh04, Ruh05a, Ruh05b, Ruh05c, Ruh05d, Ruh06b, Ruh06c, Ruh07d, Ruh07f, Ruh08b, Ruh08c, Ruh08d, Ruh09c, Ruh09d, Ruh09b, Ruh10c, Ruh10d, Ruh11c, Ruh11d, Ruh12b, Ruh12c, Ruh07c, Ruh07e]. Convergence [Ruh67, Ruh68]. convergent [Ruh75a]. criteria [ZR86].

Data [MC87, Ruh80b, Ruh80c]. Dedication [GKRV90]. Determination [Ruh70c].

Editorial [Ruh03a, Ruh03b, Ruh06a, Ruh07a, Ruh07b, Ruh08a, Ruh10a, Ruh10b, Ruh11b, Ruh12a]. eigenproblem [Ruh70b]. Eigenvalue [Ruh73, Ruh84, Ruh93, RS98, BDDR00, ER80, OR06, Ruh70a, Ruh74a, Ruh74c, Ruh75a, Ruh76, Ruh83b, Ruh94c, Ruh94d, Ruh94b, Ruh00b, Ruh00a, ZR86]. Eigenvalues [Ruh66a, Ruh66b, Ruh70d, Ruh79a, Ruh80a, Ruh75b, Ruh77, Ruh90, Ruh96, Ruh10e]. eigenvectors [Ruh77]. Elementary [Ruh80d]. elimination [Ruh79b]. empirical [Ruh80b, Ruh80c]. environment [Ruh80a]. equality [Ruh79b]. Erik [Ruh94a]. example [HR90]. exponentials [Ruh80b, Ruh80c].

F2 [KR80a]. Field [ER87]. finally [Ruh87]. Fitting [Ruh80b, Ruh80c]. flow [HR90]. following [HR90, ZR86]. Form [KR80a, KR80b]. formation [AHR87]. found [Ruh87]. Fröberg [Ruh94a]. Functions [MC87, RK78].

Gauss [Ruh79c]. Gene [Ruh08a]. General [Ruh70c]. Generalization [Ruh68]. Generalized [Ruh93, ER80, Ruh79b, WRWD84]. Golub [Ruh08a]. gradients [RW72]. Gram [Ruh83a]. guide [BDDR00].

Havsbad [KR83]. Held [KR83, MC87]. her [GKRV90]. Howard [Ruh08a]. hydrodynamical [HR90].

II [Ruh94d]. III [Ruh94b]. ill [Ruh70b]. ill-conditioned [Ruh70b]. IMA [MC87]. Implementation [Ruh79a]. Information [BR01, BR05]. Introduction [Ruh04, Ruh05a, Ruh05b, Ruh05c, Ruh05d, Ruh06b, Ruh06c, Ruh07c, Ruh07d, Ruh07e, Ruh07f, Ruh08b, Ruh08c, Ruh08d, Ruh09b, Ruh09c, Ruh09d, Ruh10c, Ruh10d, Ruh11c, Ruh11d, Ruh12b, Ruh12c]. Invariant [Ruh70d]. Inverse [Ruh80d, RW72]. inverses [WRWD84]. Issue [Ruh04, Ruh05a, Ruh05b, Ruh05c, Ruh05d, Ruh06b, Ruh06c, Ruh07d, Ruh07f, Ruh08b, Ruh08c, Ruh08d, Ruh09c, Ruh09d, Ruh09b, Ruh10c, Ruh10d, Ruh11c, Ruh11d, Ruh12b, Ruh12c, Ruh07c, Ruh07e]. Iteration [Ruh80a, Ruh80d, RW72]. Iterative [Ruh74a, Ruh75a, GGL94].

Jacobi [Ruh67, Ruh68, Ruh80d]. JNF [KR80a]. Jordan [KR80a, KR80b]. July [MC87]. June [Ruh94a].

Krylov [BR01, BR05, OR06, Ruh84, Ruh93, Ruh94c, Ruh94d, Ruh94b, RS98, Ruh98, Ruh00b, Ruh00a, Ruh10e]. Kublanovskaya [GKRV90].

Lanczos [ER80, ER82, ER87, Ruh79a]. Large [Ruh79a, Ruh98, ER80, Ruh74a, Ruh74c, Ruh76]. Least [Ruh79c, RW80, Ruh79b, WRWD84]. Linear [BDD^{+}95, Ruh82, Ruh79b, WRWD84].

March [KR83]. Mathematical [Ruh82]. Matrices [Ruh67, Ruh68, Ruh79a, Ruh74a, Ruh74c, Ruh75b, Ruh94b]. Matrix [ER87, KR80a, KR80b, Ruh69, Ruh70c, RK78, Ruh98, Ruh66a, Ruh66b, Ruh70b, Ruh87, Ruh95, Ruh00b, Ruh00a, KR83, Ruh94d]. Means [Ruh70d]. memoriam [Ruh08a]. Method [Ruh67, Ruh68, BR05, ER80, Ruh66a, Ruh66b, RW72]. Methods [Ruh84, GGL94, Ruh74c, Ruh76]. Military [MC87]. missing [Ruh74b]. Model [RS98, OR06].

N [GKRV90]. Newton [Ruh79c]. Nonlinear [Ruh73, Ruh79c, RW80, Ruh96, Ruh00b, Ruh00a]. Nonsymmetric [Ruh98, Ruh70a, Ruh83b, Ruh94c, Ruh94d, Ruh94b]. Norm [Ruh69]. Normal [KR80a, KR80b, Ruh67, Ruh75b, Ruh87]. Numerical [KR80a, KR80b, Ruh70c, Ruh74b, Ruh76, RK78, Ruh83a, ZR86, ER80, HR90].

observations [Ruh74b]. order [OR06]. Orthogonalization [Ruh83a].

package [ER82]. pairs [Ruh94d]. partial [WRWD84]. path [HR90, ZR86]. Pencils [ER87, KR83, Ruh98, Ruh10e]. Perturbation [Ruh70d]. Pite [KR83]. PLS [WRWD84]. Polynomial [Ruh95]. positive [Ruh80b, Ruh80c]. Practical [Ruh98, BDDR00]. Preface [BBRZ02, Ruh06d, Ruh07f, Ruh09d, Ruh12d, Ruh13a, Ruh13b, Ruh13c, Ruh13d, Ruh14a, Ruh14b, Ruh14c]. principal [Ruh74b]. Problem [Ruh73, AHR87, Ruh70a, Ruh74c, WRWD84]. Problems [BDD^{+}95, Ruh79c, RW80, BDDR00, ER80, Ruh76, Ruh79b, Ruh83b, Ruh93, Ruh94c, Ruh94d, Ruh94b, Ruh00b, Ruh00a]. Proceedings [KR83, MC87]. programming [Ruh66a]. Properties [Ruh70b].

Quadratic [Ruh67, Ruh68]. Quotient [Ruh80a].

Rational [OR06, Ruh84, Ruh93, Ruh94c, Ruh94d, RS98, Ruh98, Ruh10e, Ruh94b, Ruh00b, Ruh00a]. Rayleigh [Ruh80a]. real [Ruh94b, Ruh10e]. Reduction [RS98, OR06]. Reflections [Ruh80d]. regression [WRWD84]. Relation [Ruh80d]. retrieval [BR01, BR05]. Review [Ruh95]. Riemann [Ruh90]. Rotations [ER87]. Royal [MC87]. Ruhe [BE15].

scheme [Ruh79b]. Schmidt [Ruh83a]. Science [MC87]. Separable [RW80]. Sequence [Ruh84]. sequences [BR01]. several [Ruh74b]. shifts [Ruh94b]. short [BR01]. Showing [Ruh90]. Shrivenham [MC87]. sided [Ruh83b]. Similarity [Ruh69]. singular [Ruh75b]. Software [Ruh82, ER82]. solution [AHR87, BDDR00, ER80, Ruh76]. SOR [Ruh74c]. SOR-methods [Ruh74c]. Sparse [Ruh79a, Ruh98, ER80, Ruh74c, Ruh76]. Spectral [ER82, ER80, Ruh96]. sphere [Ruh90]. splittings [Ruh75a]. Squares [Ruh79c, RW80, Ruh79b, WRWD84]. STLM [ER82]. Structure [Ruh70c]. subspace [BR05]. Subspaces [Ruh70d]. sums [Ruh80b, Ruh80c]. Sweden [KR83]. switching [ZR86]. Symmetric [ER87, Ruh79a, ER80, Ruh74a].

Templates [BDD^{+}95, BDDR00]. Transformation [ER80, ER82, Ruh69, Ruh96]. two [Ruh83b]. two-sided [Ruh83b].

universal [Ruh79b]. used [RW72]. using [BR01].

Value [ER87]. values [Ruh75b]. Vectors [Ruh83a]. Vera [GKRV90]. very [BR01, Ruh70b]. volume [Ruh09a].

References

Andersson:1987:CFC

[AHR87] L. Andersson, K. Holmström, and A. Ruhe. Complex formation constants — a problem from solution chemistry. In Mason and Cox [MC87], pages 557–572. ISBN 0-19-853612-7. LCCN QA221 .A5361 1987; QA221 .I47 1985. US$90.

Barlow:2002:P

[BBRZ02] Jesse L. Barlow, Michael W.

Berry, Axel Ruhe, and Hongyuan Zha. Preface. Computational Statistics & Data Analysis, 41(1): 1–2, November 28, 2002. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0167947302002049.

Bai:1995:TLAa

[BDD^{+}95] Z. Bai, D. Day, J. Demmel, J. Dongarra, M. Gu, A. Ruhe, and H. van der Vorst. Templates for linear algebra problems. LAPACK Working Note 106, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October 1995.

URL http://www.netlib.org/
lapack/lawns/lawn106.ps;
http://www.netlib.org/lapack/
lawnspdf/lawn106.pdf. UT-CS-

95-311, October 1995.

Bai:2000:TSA

[BDDR00] Zhaojun Bai, James Demmel, Jack Dongarra, and Axel Ruhe, editors. Templates for the solution of algebraic eigenvalue problems: a practical guide. Software, Environments, and Tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2000. ISBN 0-89871-471-0 (paperback), 0-89871-958-5 (e-book). xxx + 410 pp. LCCN QA193 .T46 2000. URL

http://www.loc.gov/catdir/
enhancements/fy0726/00059507-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0726/
00059507-t.html. A practical

guide.

Bjorck:2015:AR

[BE15] Åke Björck and Lars Eldén. Axel Ruhe 1942–2015. BIT Numerical Mathematics, 55(3): 621–623, September 2015. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

link.springer.com/article/
10.1007/s10543-015-0579-4;
 http://link.springer.com/
content/pdf/10.1007/s10543-
015-0579-4.pdf.

Blom:2001:IRU

[BR01] Katarina Blom and Axel Ruhe.

Information retrieval using very short Krylov sequences. In Michael W. Berry, editor, Computational information retrieval. Proceedings of the conference CIR ’00, University of North Carolina, Raleigh, NC, USA, October 22, 2000, pages 39–52. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2001.

Blom:2005:KSM

[BR05] Katarina Blom and Axel Ruhe.

A Krylov subspace method for information retrieval. SIAM Journal on Matrix Analysis and Applications, 26(2):566– 582, 2005. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/39226.

Ericsson:1980:STL

[ER80] T. Ericsson and A. Ruhe. The spectral transformation Lanczos method for the numerical solution of large sparse generalized symmetric eigenvalue problems. Mathematics of Computation, 35: 1251–1268, 1980.

Ericsson:1982:SSP

[ER82] Thomas Ericsson and Axel Ruhe.

STLM — a software package for the spectral transformation Lanczos algorithm. Technical report, UMINF-101, 1982. 1–32 pp.

Ericsson:1987:LAF

[ER87] Thomas Ericsson and Axel Ruhe. Lanczos algorithms and field of value rotations for symmetric matrix pencils. Linear Algebra and its Applications, 88/89:733– 746, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Golub:1994:RAI

[GGL94] Gene Golub, Anne Greenbaum, and Mitchell Luskin, editors. Recent advances in iterative methods, volume 60 of The IMA Volumes in Mathematics and its Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-94252-1. LCCN QA297.8 .R43 1994.

Golub:1990:DVK

[GKRV90] Gene Golub, Bo Kågström, Axel Ruhe, and Paul Van Dooren. Dedication to Vera N. Kublanovskaya on her 70th birthday. SIAM Journal on Matrix Analysis and Applications, 11 (4):vii–??, October 1990. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Huitfeldt:1990:NAN

[HR90] Jacques Huitfeldt and Axel Ruhe. A new algorithm for numerical path following applied to an example from hydrodynamical flow. SIAM J. Sci. Stat. Comp., 11(6): 1181–1192, 1990. CODEN SI-JCD4. ISSN 0196-5204.

Kaagstroem:1980:AJA

[KR80a] Bo Kågström and Axel Ruhe. Algorithm 560: JNF, an algorithm for numerical computation of the Jordan normal form of a complex matrix [F2]. ACM Transactions on Mathematical Software, 6(3): 437–443, September 1980. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kaagstroem:1980:ANC

[KR80b] Bo Kågström and Axel Ruhe. An algorithm for numerical computation of the Jordan normal form of a complex matrix. ACM Transactions on Mathematical Software, 6(3):398–419, September 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kaagstrom:1983:MPP

[KR83] Bo Kågström and Axel Ruhe, editors. Matrix Pencils: Proceedings of a Conference Held at Pite Havsbad, Sweden, March 22–24, 1982, volume 973 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CO-DEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 973. URL http:/

/link.springer.com/book/10.
1007/BFb0062089;   http://
www.springerlink.com/content/
978-3-540-39447-1.

Mason:1987:AAB

[MC87] J. C. Mason and M. G. Cox, editors. Algorithms for approximation: based on the proceedings of the IMA Conference on Algorithms for the Approximation of Functions and Data, held at the Royal Military College of Science, Shrivenham, July 1985, volume 10 of The Institute of Mathematics and Its Applications conference series, new series. Clarendon Press, Oxford, UK, 1987. ISBN 0-19-853612-7. LCCN QA221 .A5361 1987; QA221 .I47 1985. US$90.

Olsson:2006:RKE

[OR06] K. Henrik A. Olsson and Axel Ruhe. Rational Krylov for eigenvalue computation and model order reduction. BIT Numerical Mathematics, 46(S-1):99–111, November 2006. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1978:NCM

[RK78] A. Ruhe and B. Kågstrom. Numerical computation of matrix functions. SIAM Review, 20 (3):634, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruhe:1998:RKA

[RS98] Axel Ruhe and Daniel Skoogh. Rational Krylov algorithms for eigenvalue computation and model reduction. Lecture Notes in Computer Science, 1541:491– 502, 1998. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruhe:1966:APE

[Ruh66a] Axel Ruhe. Algol programming:

Eigenvalues of a complex matrix by the QR-method. Nordisk tidskrift for informationsbehandling, 6(4):350–358, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=6&issue=4&spage=350.

Ruhe:1966:ECM

[Ruh66b] Axel Ruhe. Eigenvalues of a complex matrix by the QRmethod. Nordisk tidskrift for informationsbehandling, 6(??):350– 358, 1966.

Ruhe:1967:QCJ

[Ruh67] Axel Ruhe. On the quadratic convergence of the Jacobi method for normal matrices. BIT, 7(4): 305–313, 1967. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1968:QCG

[Ruh68] Axel Ruhe. On the quadratic convergence of a generalization of the Jacobi method to arbitrary matrices. BIT, 8(3): 210–231, 1968. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1969:NMA

[Ruh69] Axel Ruhe. The norm of a matrix after a similarity transformation.

BIT, 9(1):53–58, 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1970:ANA

[Ruh70a] A. Ruhe. Algorithms for the nonsymmetric algebraic eigenvalue problem. Ph.D. thesis, Lund University, Lund, Sweden, 1970.

Ruhe:1970:PMV

[Ruh70b] A. Ruhe. Properties of a matrix with a very ill-conditioned eigenproblem. Num. Math, 15: 57–60, 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ruhe:1970:AND

[Ruh70c] Axel Ruhe. An algorithm for numerical determination of the structure of a general matrix. BIT, 10(2):196–216, 1970. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1970:PBM

[Ruh70d] Axel Ruhe. Perturbation bounds for means of eigenvalues and invariant subspaces. BIT, 10(3): 343–354, 1970. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1973:ANE

[Ruh73]  Axel Ruhe.  Algorithms for

the nonlinear eigenvalue problem. SIAM Journal on Numerical Analysis, 10(4):674–689, September 1973. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Ruhe:1974:IEA

[Ruh74a] A. Ruhe. Iterative eigenvalue algorithms for large symmetric matrices. In L. Collatz and K. P. Hadeler, editors, Eigenwertprobleme, ISNM 24, volume 24, pages 97–115. Birkhäuser Verlag, Basel, 1974.

Ruhe:1974:NCP

[Ruh74b] A. Ruhe. Numerical computation of principal components when several observations are missing. Technical report, UMINF-48, Umeå, Sweden, 1974.

Ruhe:1974:SME

[Ruh74c] A. Ruhe. SOR-methods for the eigenvalue problem of large sparse matrices. Mathematics of Computation, 28:695–710, 1974.

Ruhe:1975:IEA

[Ruh75a] A. Ruhe. Iterative eigenvalue algorithms based on convergent splittings. J. Computational Physics, 19:110–120, 1975.

Ruhe:1975:CES

[Ruh75b] A. Ruhe. On the closeness of eigenvalues and singular values for almost normal matrices. Linear Algebra and its Applications, 11(1):87–93, 1975. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379575901196.

Ruhe:1976:NMS

[Ruh76] A. Ruhe. Numerical methods for the solution of large sparse eigenvalue problems. In V. A. Barker, editor, Sparse Matrix Techniques, Lect Notes Math 572, pages 130– 184. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976.

Ruhe:1977:CEE

[Ruh77] Axel Ruhe. Computation of eigenvalues and eigenvectors. Lecture Notes in Mathematics, 572: 130–184, 1977. CODEN LN-MAA2. ISBN 3-540-08130-5 (print), 3-540-37430-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0116617/.

Ruhe:1979:IAB

[Ruh79a] A. Ruhe. Implementation aspects of band Lanczos algorithms for computation of eigenvalues of large sparse symmetric matrices. Mathematics of Computation, 33: 680–687, 1979.

Ruhe:1979:UES

[Ruh79b] A. Ruhe. A universal elimination scheme for generalized linear least squares problems with equality constraints. Technical report, UMINF-74, 1979.

Ruhe:1979:AGN

[Ruh79c] Axel Ruhe. Accelerated Gauss– Newton algorithms for nonlinear least squares problems. BIT, 19 (3):356–367, 1979. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1980:EAE

[Ruh80a] A. Ruhe. Eigenvalues in an APL environment, an algorithm based on Rayleigh quotient iteration. APL Quote Quad, 10(3): 29–30, 1980. CODEN APLQD9. ISSN 0163-6006 (print), 1558-3392 (electronic).

Ruhe:1980:FEDa

[Ruh80b] A. Ruhe. Fitting empirical data by positive sums of exponentials. SIAM J. Sci. Stat. Comp., 1: 481–498, 1980. CODEN SIJCD4. ISSN 0196-5204.

Ruhe:1980:FEDb

[Ruh80c] A. Ruhe. Fitting empirical data by positive sums of exponentials. In M. M. Barritt and D. Wishart, editors, COMPSTAT 1980: proceedings in computational statistics; 4th symposium held at Edinburgh 1980, pages 622–628. Physica-Verlag, Vienna, Austria, 1980. ISBN 3-7908-0229-8. LCCN QA276.4 .C195.1980.

Ruhe:1980:RBJ

[Ruh80d] Axel Ruhe. The relation between the Jacobi algorithm and inverse iteration and a Jacobi algorithm based on elementary reflections. BIT, 20(1):88–96, 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1982:MSL

[Ruh82] A. Ruhe. Mathematical software for linear algebra computations. Technical report, UMINF-95, 1982. 1–15 pp.

Ruhe:1983:NAG

[Ruh83a] Axel Ruhe. Numerical aspects of Gram–Schmidt orthogonalization of vectors. Linear Algebra and its Applications, 52/53: 591–601, 1983. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ruhe:1983:TSA

[Ruh83b] Axel Ruhe. The two-sided Arnoldi algorithm for nonsymmetric eigenvalue problems. Lecture Notes in Mathematics, 973: 104–120, 1983. CODEN LN-MAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

                                        link.springer.com/chapter/
                                        10.1007/BFb0062097/.

Ruhe:1984:RKS

[Ruh84] Axel Ruhe. Rational Krylov sequence methods for eigenvalue computation. Linear Algebra and its Applications, 58:391– 405, 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ruhe:1987:CNM

[Ruh87] Axel Ruhe. Closest normal matrix finally found! BIT, 27(4):

585–598, 1987. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1990:SER

[Ruh90]  Axel Ruhe.  Showing eigen-

values on the Riemann sphere. In Bo Kågström, editor, Visualization and analysis of scientific and technical computations on high-performance computers, pages 136–141. University of Umeå, Report UMINF-191.90, 1990.

Ruhe:1993:RKA

[Ruh93] Axel Ruhe. The rational Krylov algorithm for generalized eigenvalue problems. In 92 Shanghai International Numerical Algebra and its applications Conference, Fudan University, Shanghai, October 26–30, 1992, to appear. ????, ????, 1993.

Ruhe:1994:CEF

[Ruh94a] Axel Ruhe. Carl-Erik Fröberg on his 75th birthday June 23, 1993. BIT, 34(1):i–ii, March 1994. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/
BIT/contents/bit34.html;
http://www.springerlink.com/
openurl.asp?genre=article&
issn=0006-3835&volume=34&issue=
1&spage=??.

Ruhe:1994:RKAc

[Ruh94b] Axel Ruhe. The rational Krylov algorithm for nonsymmetric eigenvalue problems. III:

Complex shifts for real matrices. BIT, 34(1):165–176, March 1994. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

                                        URL http://www.mai.liu.se/
                                        BIT/contents/bit34.html.

Ruhe:1994:RKAa

[Ruh94c] Axel Ruhe. Rational Krylov algorithms for nonsymmetric eigenvalue problems. In Golub et al. [GGL94], pages 149–164. ISBN 0-387-94252-1. LCCN QA297.8 .R43 1994.

Ruhe:1994:RKAb

[Ruh94d] Axel Ruhe. Rational Krylov algorithms for nonsymmetric eigenvalue problems. II. Matrix pairs. Linear Algebra and its Applications, 197/198:283–295, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Ruhe:1995:BRP

[Ruh95] Axel Ruhe. Book review: Polynomial and matrix computations. Bulletin of the American Mathematical Society (new series), 32(3):357–360, 1995. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Ruhe:1996:CNE

[Ruh96] Axel Ruhe. Computing nonlinear eigenvalues with spectral transformation Arnoldi. Zeitschrift für Angewandte Mathematik und Mechanik, 76:17–20, 1996. CO-DEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Ruhe:1998:RKP

[Ruh98] Axel Ruhe. Rational Krylov:

a practical algorithm for large sparse nonsymmetric matrix pencils. SIAM Journal on Scientific Computing, 19(5):1535– 1551, September 1998. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28559.

Ruhe:2000:RKAb

[Ruh00a] A. Ruhe. The rational Krylov algorithm for nonlinear matrix eigenvalue problems. Journal of Mathematical Sciences (New York), 114(6):1854–1856, 2000. CODEN JMTSEW. ISSN 1072-3374 (print), 1573-8795 (electronic). URL http:/

/link.springer.com/article/
10.1023/A:1022414805217.

Ruhe:2000:RKAa

[Ruh00b] Axel Ruhe. A rational Krylov algorithm for nonlinear matrix eigenvalue problems. Zap. Nauchn. Sem. S.-Peterburg. Otdel. Mat. Inst. Steklov. (POMI), 268(Chisl. Metody i Vopr. Organ. Vychisl. 14):176–180, 245– 246, 2000. ISSN 0373-2703. English translation in Journal of Mathematical Sciences, 114(6), 1854–1856 (April 2003).

Ruhe:2003:Ea

[Ruh03a] Axel Ruhe. Editorial. BIT Numerical Mathematics, 43(1): v–viii, March 2003. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=43&issue=1&spage=??.

Ruhe:2003:Eb

[Ruh03b] Axel Ruhe. Editorial. BIT Numerical Mathematics, 43(2): iii, June 2003. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=43&issue=2&spage=??.

Ruhe:2004:ICI

[Ruh04] Axel Ruhe. Introduction to the contents of issue 44:4. BIT Numerical Mathematics, 44(4): 5, December 2004. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=44&issue=4&spage=5.

Ruhe:2005:ICIa

[Ruh05a] Axel Ruhe. Introduction to the contents of issue 45:1. BIT Numerical Mathematics, 45(1): ??, March 2005. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=45&issue=1&spage=??.

Ruhe:2005:ICIb

[Ruh05b] Axel Ruhe. Introduction to the contents of issue 45:2. BIT Numerical Mathematics, 45(2): ??, June 2005. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=45&issue=2&spage=??.

Ruhe:2005:ICIc

[Ruh05c] Axel Ruhe. Introduction to the contents of issue 45:3. BIT Numerical Mathematics, 45(3):??, September 2005. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=45&issue=3&spage=??.

Ruhe:2005:ICId

[Ruh05d] Axel Ruhe. Introduction to the contents of issue 45:4. BIT Numerical Mathematics, 45(4):vii, December 2005. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=45&issue=4&spage=??.

Ruhe:2006:E

[Ruh06a] Axel Ruhe. Editorial. BIT Numerical Mathematics, 46(4):685– 686, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:2006:ICIa

[Ruh06b] Axel Ruhe. Introduction to the contents of issue 46:1. BIT Numerical Mathematics, 46(1): 3, March 2006. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:2006:ICIb

[Ruh06c] Axel Ruhe. Introduction to the contents of issue 46:4. BIT Numerical Mathematics, 46(4):687– 688, December 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:2006:P

[Ruh06d] Axel Ruhe. Preface. BIT Numerical Mathematics, 46(1): 1, March 2006. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:2007:Ea

[Ruh07a] Axel Ruhe. Editorial. BIT Numerical Mathematics, 47(3): 483, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/openurl.asp?
                                        genre=article&issn=0006-3835&
                                        volume=47&issue=3&spage=483.

Ruhe:2007:Eb

[Ruh07b] Axel Ruhe. Editorial. BIT Numerical Mathematics, 47(4):693– 695, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.

springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=47&issue=4&spage=693.

Ruhe:2007:ICIa

[Ruh07c] Axel Ruhe. Introduction to the Contents of Issue 47:2. BIT Numerical Mathematics, 47(2): 239–240, June 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=47&issue=2&spage=239.

Ruhe:2007:ICIb

[Ruh07d] Axel Ruhe. Introduction to the contents of issue 47:3. BIT Numerical Mathematics, 47(3):485– 486, September 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=47&issue=3&spage=485.

Ruhe:2007:ICIc

[Ruh07e] Axel Ruhe. Introduction to the Contents of Issue 47:4. BIT Numerical Mathematics, 47(4): 697, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=47&issue=4&spage=697.

Ruhe:2007:PIC

[Ruh07f] Axel Ruhe. Preface and introduction to the contents of issue 47:1.

BIT Numerical Mathematics, 47 (1):1–2, March 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=47&issue=1&spage=1.

Ruhe:2008:EGH

[Ruh08a] Axel Ruhe. Editorial: Gene Howard Golub in memoriam. BIT Numerical Mathematics, 48(1):1– 2, March 2008. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=48&issue=1&spage=1.

Ruhe:2008:ICIa

[Ruh08b] Axel Ruhe. Introduction to the contents of issue 48:1. BIT Numerical Mathematics, 48(1): 3, March 2008. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=48&issue=1&spage=3.

Ruhe:2008:ICIb

[Ruh08c] Axel Ruhe. Introduction to the contents of issue 48:3. BIT Numerical Mathematics, 48(3):423– 424, September 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?

genre=article&issn=0006-3835&
volume=48&issue=3&spage=423.

Ruhe:2008:ICIc

[Ruh08d] Axel Ruhe. Introduction to the contents of issue 48:4. BIT Numerical Mathematics, 48(4):649– 650, December 2008. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=48&issue=4&spage=649.

Ruhe:2009:CV

[Ruh09a] Axel Ruhe. Collaborators for volume 49. BIT Numerical Mathematics, 49(4):641–643, December 2009. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=49&issue=4&spage=641.

Ruhe:2009:ICIa

[Ruh09b] Axel Ruhe. Introduction to the contents of issue 49:2. BIT Numerical Mathematics, 49(2):247– 248, June 2009. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=49&issue=2&spage=247.

Ruhe:2009:ICIb

[Ruh09c] Axel Ruhe. Introduction to the contents of issue 49:3. BIT Numerical Mathematics, 49(3):475– 476, September 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=49&issue=3&spage=475.

Ruhe:2009:PIC

[Ruh09d] Axel Ruhe. Preface and introduction to the contents of issue 49:1. BIT Numerical Mathematics, 49 (1):1–2, March 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=49&issue=1&spage=1.

Ruhe:2010:Ea

[Ruh10a] Axel Ruhe. Editorial. BIT Numerical Mathematics, 50(3):451– 453, September 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=50&issue=3&spage=451.

Ruhe:2010:Eb

[Ruh10b] Axel Ruhe. Editorial. BIT Numerical Mathematics, 50(4):689– 692, December 2010. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=50&issue=4&spage=689.

Ruhe:2010:ICIa

[Ruh10c] Axel Ruhe. Introduction to the contents of issue 50:1. BIT Numerical Mathematics, 50(1):1– 2, March 2010. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=50&issue=1&spage=1.

Ruhe:2010:ICIb

[Ruh10d] Axel Ruhe. Introduction to the contents of issue 50:2. BIT Numerical Mathematics, 50(2):221– 222, June 2010. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=50&issue=2&spage=221.

Ruhe:2010:RKR

[Ruh10e] Axel Ruhe. Rational Krylov for real pencils with complex eigenvalues. Taiwanese Journal of Mathematics, 14(3A):795–803, 2010. ISSN 1027-5487.

Ruhe:2011:B

[Ruh11a] Axel Ruhe. BIT50. BIT Numerical Mathematics, 51(1):1– 5, March 2011. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=51&issue=1&spage=1.

Ruhe:2011:E

[Ruh11b] Axel Ruhe. Editorial. BIT Numerical Mathematics, 51(4):777– 780, December 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=51&issue=4&spage=777.

Ruhe:2011:ICIa

[Ruh11c] Axel Ruhe. Introduction to the contents of issue 51:2. BIT Numerical Mathematics, 51(2):251– 252, June 2011. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=51&issue=2&spage=251.

Ruhe:2011:ICIb

[Ruh11d] Axel Ruhe. Introduction to the contents of issue 51:3. BIT Numerical Mathematics, 51(3):481– 482, September 2011. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=51&issue=3&spage=481.

Ruhe:2012:E

[Ruh12a] Axel Ruhe. Editorial. BIT Numerical Mathematics, 52(4): 797–800, December 2012. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

link.springer.com/article/
10.1007/s10543-012-0408-y;
 http://link.springer.com/
content/pdf/10.1007/s10543-
012-0408-y.

Ruhe:2012:ICIc

[Ruh12b] Axel Ruhe. Introduction to the contents of issue 52:2. BIT Numerical Mathematics, 52(2):271– 272, June 2012. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=52&issue=2&spage=271.

Ruhe:2012:ICId

[Ruh12c] Axel Ruhe. Introduction to the contents of issue 52:3. BIT Numerical Mathematics, 52(3):537– 538, September 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=52&issue=3&spage=537.

Ruhe:2012:PB

[Ruh12d] Axel Ruhe. Preface to BIT 52:1.

BIT Numerical Mathematics, 52 (1):1–2, March 2012. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=52&issue=1&spage=1.

Ruhe:2013:PBa

[Ruh13a] Axel Ruhe. Preface to BIT 53:1. BIT Numerical Mathematics, 53(1):1–2, March 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:/

                                        /link.springer.com/article/
                                        10.1007/s10543-013-0427-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10543-
                                        013-0427-3.

Ruhe:2013:PBb

[Ruh13b] Axel Ruhe. Preface to BIT 53:2. BIT Numerical Mathematics, 53(2):283–284, June 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:/

                                        /link.springer.com/article/
                                        10.1007/s10543-013-0428-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10543-
                                        013-0428-2.pdf.

Ruhe:2013:PBc

[Ruh13c] Axel Ruhe. Preface to BIT 53:3. BIT Numerical Mathematics, 53 (3):565–566, September 2013. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:/

                                        /link.springer.com/article/
                                        10.1007/s10543-013-0440-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10543-
                                        013-0440-6.pdf.

Ruhe:2013:PBd

[Ruh13d] Axel Ruhe. Preface to BIT 53:4. BIT Numerical Mathematics, 53(4):821–825, December 2013.

CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:/

/link.springer.com/article/
10.1007/s10543-013-0456-y;
 http://link.springer.com/
content/pdf/10.1007/s10543-
013-0456-y.pdf.

Ruhe:2014:P

[Ruh14a] Axel Ruhe. Preface to 54-3. BIT Numerical Mathematics, 54(3):585–586, September 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:/

/link.springer.com/article/
10.1007/s10543-014-0518-9;
 http://link.springer.com/
content/pdf/10.1007/s10543-
014-0518-9.pdf.

Ruhe:2014:PBa

[Ruh14b] Axel Ruhe. Preface to BIT 54:2. BIT Numerical Mathematics, 54(2):303–304, June 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:/

/link.springer.com/article/
10.1007/s10543-014-0504-2;
 http://link.springer.com/
content/pdf/10.1007/s10543-
014-0504-2.pdf.

Ruhe:2014:PBb

[Ruh14c] Axel Ruhe. Preface to BIT 54:4.

BIT Numerical Mathematics, 54(4):867–871, December 2014. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:/

/link.springer.com/article/

                                        10.1007/s10543-014-0535-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10543-
                                        014-0535-8.pdf.

Ruhe:1972:MCG

[RW72] Axel Ruhe and T. Wiberg. The method of conjugate gradients used in inverse iteration. BIT, 12 (4):543–554, 1972. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1980:ASN

[RW80] Axel Ruhe and Per Åke Wedin. Algorithms for separable nonlinear least squares problems. SIAM Review, 22(3):318–337, July 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wold:1984:CPL

[WRWD84] S. Wold, A. Ruhe, H. Wold, and W. J. Dunn. The collinearity problem in linear regression. the partial least squares (PLS) approach to generalized inverses. SIAM J. Sci. Stat. Comp., 5: 735–743, 1984. CODEN SIJCD4. ISSN 0196-5204.

Zhou:1986:NPF

[ZR86] Yong Feng Zhou and Axel Ruhe. Numerical path following and eigenvalue criteria for branch switching. In J. Cullum and R. A. Willoughby, editors, Large Scale Eigenvalue Problems, pages 121– 142. Elsevier Science Publishers (North-Holland), New York, NY, USA, 1986.